### v1

@@ Abstract
[SUMMARY] The attention mechanism has been proven to be an effective way to improve spiking neural network (SNN). However, based on the fact that the current SNN input data flow is split into tensors to process on GPUs, none of the previous works consider the properties of tensors to implement an attention module. This inspires us to rethink current SNN from the perspective of tensor-relevant theories. Using tensor decomposition, we design the projected full attention (PFA) module, which demonstrates excellent results with linearly growing parameters. Specifically, PFA is composed by the linear projection of spike tensor (LPST) module and attention map composing (AMC) module. In LPST, we start by compressing the original spike tensor into three projected tensors using a single property-preserving strategy with learnable parameters for each dimension. Then, in AMC, we exploit the inverse procedure of the tensor decomposition process to combine the three tensors into the attention map using a so-called connecting factor. To validate the effectiveness of the proposed PFA module, we integrate it into the widely used VGG and ResNet architectures for classification tasks. Our method achieves state-of-the-art performance on both static and dynamic benchmark datasets, surpassing the existing SNN models with Transformer-based and CNN-based backbones.
@@ 1 Introduction
Spiking neural network (SNN) is currently attracting the interest of academics due to its lower energy consumption and greater bio-interpretability compared to traditional artificial neural network (ANN). The brain-inspired computation model gives it significant potential in processing temporal data. Though training a deep SNN is still a challenge, recent developments in SNN introduce backpropagation [3] to relieve this issue to some degree. Moreover, it makes the extension of ANN modules into SNN possible, such as batch normalization and residual blocks.
@@
As it is now feasible to incorporate ANN modules into SNNs while maintaining their inherent power efficiency, we can leverage these modules to augment the overall performance of SNNs. For instance, Zheng et al. [45] propose threshold-dependent batch normalization to solve the problems of gradient vanishing and threshold-input balance. The well-known residual block proposed by He et al. [14] makes it possible to train an ultra-deep ANN network. To introduce residual block into the field of SNN, Hu et al. [17] investigate converting residual networks and propose a new technique to adjust the parameter weight correspondingly. Furthermore, Fang et al. [11] propose the SEW-ResNet to solve gradient vanishing problems and achieve identity mapping. In addition, Hu et al. [18] design several novel residual blocks and test their applicability, greatly extending the depth of a directly trained SNN. In summary, several strategies in ANN, such as normalization, architecture, and training methods, could favor the performance of SNN effectively, which also motivates us to explore more useful techniques in ANNs for better-conducting SNNs.
@@
Figure 1: Accuracy on CIFAR10 (left) and CIFAR100 (right). Compared with other methods, PFA significantly improves the network performance.
@@
As one of the most promising techniques in ANNs, the attention mechanism can mimic the human ability to focus on certain things while ignoring others. Recently, three representative works have proven that the attention mechanism can successfully work in SNNs. Especially, Yao et al. [42] switch channel attention to the temporal dimension, revealing that paying attention to the temporal dimension is effective for performance enhancement. Zhu et al. [48] apply two 1D convolutions to a generated 2D tensor to achieve temporal-channel attention. Inspired by experimental observations on predictive attentional remapping, Cai et al. [4] design the SCTFA module to assess the input tensor from temporal, channel, and spatial dimensions, achieving attention among these three dimensions.
@@
Current software frameworks, i.e. PyTorch, and tensorflow, are boosted by GPU using parallel computation. This makes it hard to realize event-driven as events occur asynchronously. Based on these frameworks, as a trade-off, a popular approach is splitting the input event streams into slices and composing tensors to convenient later computation. This approach is adopted by all attention mechanism-related works, showing a promising outcome. Since we can now only efficiently process event-converted tensors, it is natural for us to consider involving tensor analysis theories to implement an attention module for SNN, which is ignored by previous work. Simultaneously, this analysis highlights the inherent resilience within SNN modules. It demonstrates that even when employing low-rank execution for Attention mechanisms, substantial performance enhancements can be achieved. This aligns perfectly with the sparsity philosophy of SNNs.
In this paper, we propose the projected-full attention (PFA), which considers projecting one input tensor into three separate lower-dimensional tensors, then calculate a full-attention map by utilizing the inverse procedure of the tensor decomposition, fully exploiting the peculiar temporal dimension in SNN with minimal computing complexity. The input tensor is first projected along three separate dimensions by the linear projection of spike tensor (LPST) module. Then, the projections are composed into an attention map by the attention map composing (AMC) module. The input tensor is fused with the attention map by the Hadamard product to obtain the refined tensor. The main contributions of this paper can be summarized as follows:
  • We propose the projected full-attention (PFA) module to achieve three-dimensional attention. Compared with previous works, PFA is a lightweight module with fewer learnable parameters, and the number of parameters grows linearly as the size of input data increases. PFA is founded on a well-established mathematical tool, thus enhancing its reliability and mathematical interpretability.
@@
  • We present a comprehensive theoretical analysis that encompasses two key aspects: the rank of tensors and connecting factor. Based on these analyses, we propose a selection criterion for that is tailored for distinct datasets. This criterion effectively elucidates the experimental outcomes and offers valuable guidance for future applications.
@@
  • We conduct extensive experiments on both static and dynamic datasets. Experimental results show that our PFA is very effective and achieves state-of-the-art (SOTA) accuracy on both dynamic and static datasets. In addition, ablation studies demonstrate the structural reasonableness of PFA. We visualize the attention map obtained by PFA to intuitively illustrate the attention distribution among temporal-channel-spatial dimensions.
2 Related Works and Motivation
Spiking Neural Network: Spiking neural network (SNN), the third generation of artificial neural network [29], has the potential to behave as efficiently as the human brain. In an SNN, information is transmitted through spikes, which are binary events that represent the firing of a neuron. Unlike traditional ANNs, which use continuous signals, SNN neurons communicate over time using discrete spikes. Compared with ANN, SNN is more biologically realistic by adopting models from neural dynamics and is more hardware-friendly thanks to the discrete spike. Due to its temporal property, it is supposed to have better temporal expressiveness than ANN. Quantities of studies have already been conducted. For neurodynamic models, Hodgkin and Huxley proposed the Hodgkin-Huxley (H-H) model [15] in 1952. Models like Izhikevich [19] and Leaky Integrate and Fire (LIF) [25] are suggested to strike a balance between computing complexity and bio-interpretability. Since the spiking behavior is non-differentiable, training an SNN is challenging. Several methods have been explored to address this difficulty. Some techniques [38; 39; 44] are gradient-based optimization. Others, like Rueckauer et al. [36] and Diehl et al. [9], are based on converting an ANN to SNN. There also exists ANN-SNN collaborative networks [41]. Besides the conventional datasets (static datasets) such as MNIST [7], and CIFAR10 [23], SNNs also support datasets collected from event flow cameras (dynamic datasets) like NCaltech-101 [32]. The concept behind converted datasets, of which ES-ImageNet [28], is an example to transform ordinary datasets into neural forms through specific algorithms. Another way to convert ordinary data is to recapture it using dynamic vision sensor (DVS) cameras. NCaltech-101 [32] and CIFAR10DVS [26] are in this category. As aforementioned, SNNs are hardware-friendly, making hardware implementation one of the frontier areas. Related developments include TrueNorth [2], Loihi [6], and Tianjic [33] chips. Nonetheless, there are still many obstacles in applying SNNs on various tasks, especially in its performance. In this paper, we successfully increases the classification accuracy of SNNs with the proposed PFA module.
@@ 2 Related Works and Motivation
Attention Mechanism in SNN: Since Google proposed recurrent models of visual attention in 2014 [31], the attention mechanism become a potent tool to increase the performance of a neural network, which gives distinct parts of the input data varied weights. For instance, the squeeze-and-excitation network (SENet) [16] and other attention blocks have been frequently used in traditional ANNs. From this fact, it is of vital significance to introduce the attention mechanism into SNNs. To achieve temporal-wise attention, inspired by SENet, TA-SNN [42] is proposed. Besides, TCJA [48] use two 1D convolutions to pay attention to both channel dimension and temporal dimension to prolong the attention mechanism. Nevertheless, they fail to considered three crucial aspects of SNN, i.e., temporal, channel, and spatial. By changing SENet and leaving affection in the temporal direction, Cai et al. [4] and Yao et al. [43] successfully apply temporal-channel-spatial attention to SNN. These researches reveal the fact that attention mechanism does improve the performance of SNN.
Motivation: As aforementioned, the application of the attention mechanism has demonstrated promising ability in SNN-related tasks. Regardless of the dimension(s) considered, the attention mechanism is always beneficial [42; 48; 4; 43]. Especially, the temporal-channel-spatial attention could faithfully depict the relationship of temporal, channel and spatial dimensions, thus it becomes a quite promising approach in SNNs. For this type of attention, although 3D convolution is used to accomplish a comparable effect [5], the processing complexity is intolerable. Besides, two other excellent works, i.e., Cai et al. [4] and Yao et al. [43], have fully considered the temporal-channel-spatial attention and gain exciting outcome. Nonetheless, the above approach does not take into account the nature of the input data as a tensor, which is unavoidable since the training process is accelerated by GPU using parallel computation. In this case, we naturally hope to discover properties that we can use from the relevant theories of tensors to implement the attention mechanism. Drawing inspiration from the tensor CANDECOMP/PARAFAC (CP) decomposition [22], the inverse process of the tensor CP decomposition can be used to calculate the attention map. Therefore, we propose the PFA, which mainly comprises two distinct components, i.e., LPST for the tensor compression and AMC for the attention calculation by tensor fusion. The goal of LPST is to learn parameters that help project the most important information into three 2D tensors. The AMC part uses these three projections to construct an attention map through the inverse process of CP decomposition.
3 Method
In this part, we first introduce the neuron model and the training techniques we adopt. Next, we introduce PFA and its two components, i.e. LPST and AMC, in detail. Finally, a quantitative theoretical analysis on the adoption of connecting factor $r$ is presented.
3.1 Leaky Integrate and Fire Model
In this work, we employ the Leaky Integrate and Fire (LIF) neuron due to its efficiency and simplicity. The LIF neuron accumulates external input in its membrane voltage, which can be described as follows:
$\tau\frac{dV(t)}{dt}=-(V(t)-V_{reset})+I(t),$ (1)
where $\tau$ denotes the time constant, $V(t)$ is the membrane potential at the moment $t$, and $I(t)$ is presynaptic input at $t$ moment. With this formula, the entire model is summarized as:
$\begin{cases}V_{t}^{n}=H_{t-1}^{n}+\frac{1}{\tau}[I_{t-1}^{n}-(H_{t-1}^{n}-V_{reset})]&\\ S_{t}^{n}=\Theta(V_{t}^{n}-V_{thr})\\ H_{t}^{n}=V_{t}^{n}\cdot(1-S_{t}^{n}),\end{cases}$ (2)
where $n$ and $t$ are the layer and moment indices respectively. Besides, $\tau$ represents the time constant, $V$ is membrane potential, $S$ is the spike tensor in binary form, $I$ is the input from the previous layer, $\Theta$ stands for Heaviside step function, $H$ represents the reset behavior after spiking, and $V_{thr}$ means the threshold membrane potential. The LIF model has relatively lower computation complexity and higher representational ability, which now has become a mainstream model in the field of SNN. Here, we use the spatial-temporal backpropagation technique to train our network, aiming to fully make use of the information from spatial and temporal domains. Particularly, the backpropagation on the spatial and temporal domains can be described as follows:
$\begin{cases}\frac{\partial Loss}{\partial S_{i}^{t,l}}=\sum\limits^{n^{l+1}}_{j=1}\frac{\partial Loss}{\partial S_{j}^{t,l+1}}\frac{\partial S_{j}^{t,l+1}}{\partial S_{i}^{t,l+1}}+\frac{\partial Loss}{\partial S_{j}^{t+1,l}}\frac{\partial S_{j}^{t+1,l}}{\partial S_{i}^{t,l}}\\ \frac{\partial Loss}{\partial V_{i}^{t,l}}=\frac{\partial Loss}{\partial S_{i}^{t,l}}\frac{\partial S_{i}^{t,l}}{\partial V_{i}^{t,l}}+\frac{\partial Loss}{\partial S_{i}^{t+1,l}}\frac{\partial S_{i}^{t+1,l}}{\partial V_{i}^{t,l}}.\end{cases}$ (3)
Note that the Heaviside function in (2) is non-differentiable in backpropagation. To address the issue, we employ the following approximation of the Heaviside function in backpropagation for easy calculation:
$ATan(V_{i})=\frac{1}{\pi}arctan[\frac{\pi}{2}\alpha(V_{i}-V_{thr})]+\frac{1}{2},$ (4)
where $V_{i}$ denotes the neuron membrane potential, $V_{thr}$ is the reset membrane potential and $\alpha$ is an adjustment factor to adjust the slope.

### v2

@@ Abstract
[SUMMARY] The attention mechanism has been proven to be an effective way to improve the performance of spiking neural networks (SNNs). However, from the perspective of tensor decomposition to examine the existing attention modules, we find that the rank of the attention maps generated by previous methods is fixed at 1, lacking the flexibility to adjust for specific tasks. To tackle this problem, we propose an attention module, namely Projected-full Attention (PFA), where the rank of the generated attention maps can be determined based on the characteristics of different tasks. Additionally, the parameter count of PFA grows linearly with the data scale. PFA is composed of the linear projection of spike tensor (LPST) module and attention map composing (AMC) module. In LPST, we start by compressing the original spike tensor into three projected tensors with learnable parameters for each dimension. Then, in AMC, we exploit the inverse procedure of the tensor decomposition process to combine the three tensors into the attention map using a so-called connecting factor. To validate the effectiveness of the proposed PFA module, we integrate it into the widely used VGG and ResNet architectures for classification tasks. Our method achieves state-of-the-art performance on both static and dynamic benchmark datasets, surpassing the existing SNN models with Transformer-based and CNN-based backbones. Code for PFA is available at https://github.com/RisingEntropy/PFA .
@@ 1 Introduction
Spiking neural networks (SNNs) are currently attracting the interest of academics due to their lower energy consumption and greater bio-interpretability compared to traditional artificial neural networks (ANNs)  [19; 36]. The brain-inspired computation model gives it significant potential in processing temporal data. Though training a deep SNN is still a challenge, recent developments in SNNs introduce backpropagation [3; 55; 65; 50; 53; 41] to relieve this issue to some degree. Moreover, it makes the extending of ANN modules into SNNs possible, such as batch normalization and residual blocks.
@@
As it is now feasible to incorporate ANN modules into SNNs while maintaining the inherent power efficiency, we can leverage these modules to augment the overall performance of SNNs. A good practice for this is the introduction of residual architecture to SNNs. The well-known residual block proposed by He et al. [14] makes it possible to train an ultra-deep ANN network. To introduce the residual block into the field of SNN, many variations [17; 11; 18] are presented from different perspectives, solving problems like gradient vanishing. Moreover, strategies in ANN, such as normalization [67] and architecture, could also effectively favor SNNs’ performance, thereby motivating us to explore more useful techniques in ANNs for better-conducting SNNs.
@@
Figure 1: Accuracy on CIFAR10 (left) and CIFAR100 (right). Compared with other methods, PFA significantly improves network performance.
@@
Apart from improving network architectures, numerous other approaches can improve the performance of a network. As one of the most promising techniques in ANNs, the attention mechanism can mimic the human ability to focus on certain things while ignoring others. Three representative works [62; 70; 4] have recently proven that the attention mechanism can successfully work in SNNs. Especially, Yao et al. [62] switch channel attention to the temporal dimension, revealing that paying attention to the temporal dimension is effective for performance enhancement. Zhu et al. [70] apply two 1D convolutions to a generated 2D tensor to achieve temporal-channel attention. Inspired by experimental observations on predictive attentional remapping, Cai et al. [4] design the SCTFA module to assess the input tensor from temporal, channel, and spatial dimensions, achieving attention among these three dimensions.
@@
Current software frameworks, i.e., PyTorch and Tensorflow, are boosted by GPU using parallel computation. This makes it hard to realize event-driven SNNs as events occur asynchronously. Based on these frameworks, as a trade-off, a popular approach is splitting the input event streams into slices and composing tensors for convenient later computation. This approach is adopted by all attention-mechanism-related works, showing a promising outcome. Since we are essentially dealing with tensor-formatted data flow, it is natural for us to consider involving tensor analysis theories to implement an attention module for SNNs, which is ignored by previous work. By utilizing mathematical tools such as tensor decomposition theory, we can examine attention modules from a more abstract perspective, thereby identifying the shortcomings and limitations of previous methods.
By expressing the composition of attention maps in previous methods through CANDECOMP/PARAFAC (CP) decomposition form, we found that they are essentially ‘rank-1’ methods. When applied to various tasks, they lack a certain degree of flexibility. In this paper, we introduce a module called Projected-Full Attention (PFA), which is capable of generating attention maps with a rank other than 1. Specifically, PFA consists of two components: Linear Projection of Spike Tensor (LPST) and Attention Map Composing (AMC). LPST is responsible for generating $R$ sets of projections from the input tensor using a small number of parameters. AMC is responsible for creating attention maps using these $R$ sets of projections, where $R$ is called The connecting factor representing the rank concept of the CP decomposition and is a hyper-parameter that can be adjusted based on the specific task.
The key contributions of this paper are outlined as follows:
  • We propose PFA, a module that can achieve temporal-channel-spatial attention. In contrast to previous approaches, the ranks of the attention maps generated by PFA are not fixed at 1; instead, the rank can be chosen based on the specific task. Furthermore, the parameter count of PFA grows linearly with the data scale, and the computational cost is equivalent to that of a single standard convolution layer.
@@
  • We present a comprehensive theoretical analysis, focusing on two critical aspects: the rank of tensors and $R$, the connecting factor. These analyses lead to a tailored selection criterion, providing effective insights into experimental outcomes and offering a valuable guidance for future applications.
@@
  • We conduct thorough experiments on both static and dynamic datasets. The results demonstrate the effectiveness of PFA which achieves state-of-the-art (SOTA) accuracy on both dynamic and static datasets. Ablation studies further validate the structural reasonableness of PFA. Additionally, we visualize the attention map generated by PFA to intuitively illustrate attention distribution across temporal, channel, and spatial dimensions.
2 Related Works
Spiking Neural Network: Spiking Neural Networks (SNNs), the third generation of neural networks [19; 36], offer a closer emulation of the human brain’s efficiency by utilizing discrete spikes for information transmission [64]. This unique method of communication allows for enhanced energy-saving feature compared to traditional Artificial Neural Networks (ANNs), which rely on continuous signals. The temporal representation capabilities of SNNs have been a focal point in research [62; 8; 64], drawing from seminal neurodynamic models such as the Hodgkin-Huxley (H-H) model [15] and further developed in works like Izhikevich [19] and Leaky Integrate and Fire (LIF) [31]. These models underscore SNNs’ potential in capturing time-dependent patterns in data, a feature not inherently present in ANNs. The challenges in training SNNs, due to their non-differentiable spiking behavior, have also been addressed in recent literature, with gradient-based optimization methods [54; 56; 51] and ANN-to-SNN conversion techniques [45; 9; 49; 52] being key developments. In addition to research on training methods, there are many other studies dedicated to improving the performance of SNNs, such as attention mechanisms.
@@ 2 Related Works
Attention Mechanism in SNNs: Since Google proposed recurrent models of visual attention in 2014 [38], the attention mechanism become a potent tool to increase the performance of a neural network, which gives distinct parts of the input data varied weights. For instance, the squeeze-and-excitation network (SENet) [16] gives different weights to different channels of input data and significantly improves the performance. From this fact, it is of vital significance to introduce the attention mechanism into SNNs. Qiu et al. apply the attention mechanism to the encoding layers of SNNs [42]. To achieve temporal-wise attention, inspired by SENet, TA-SNN [62] is proposed. Besides, TCJA [70] uses two 1D convolutions to pay attention to both channel dimension and temporal dimension to prolong the attention mechanism. Nevertheless, they fail to consider three crucial aspects of SNNs, i.e. temporal, channel, and spatial. By changing SENet and leaving affection in the temporal direction, Cai et al. [4] and Yao et al. [63] successfully apply temporal-channel-spatial attention to SNNs. While the methods of predecessors have been very successful, they do not take into account the fact that the input data is a high-order tensor. Examining data from a tensor perspective can provide a more mathematically abstract viewpoint, revealing additional properties. As described in Section 3.1, the methods employed by previous researchers result in ‘rank-1’ attention maps, lacking the flexibility and specificity for different tasks.
Tensor Decomposition: Tensor decomposition methods [20; 46; 25] have undergone years of development and achieved remarkable success. Many prior studies [28; 61; 39; 26; 2; 35; 48] have embraced these techniques to optimize neural network architectures or uncover versatile modules with multiple functionalities. Kossaifi et al., for instance, conducted a comprehensive review of convolution kernels from the perspective of tensor decomposition and introduced an innovative convolution module that can seamlessly extend to higher dimensions. Lau et al. propose a new decomposition method to enable the direct use of the depth-wise convolutional layer with large kernels in the attention module, without requiring any extra blocks [32]. Apart from using tensor decomposition methods to improve network structures, tensor decomposition theory can also impose certain constraints on the data itself [58; 60; 5]. Xu et al. leverage tensor decomposition to predict real-time traffic flow, demonstrating its effectiveness in handling complex patterns [60]. Chen et al. employ tensor CP decomposition to generate an attention map for segmentation, achieving significant improvements [5]. These methods either use tensor decomposition to reduce the parameter count in processing high-dimensional data or leverage low-rank properties to enhance performance. In this paper, drawing inspiration from these prior endeavors, we leverage tensor decomposition methods to enhance the field of SNNs. We design a lightweight attention module, designed to be both parameter-efficient and computationally efficient.
3 Motivation and Method
3.1 Motivation
Tensor CP decomposition [24], is a mathematical technique employed to decompose a high-order tensor into a sum of rank-one tensors. Mathematically, for a tensor denoted as $\mathcal{X}\in\mathbb{R}^{D_{1},D_{2},\dots,D_{n}}$, the CP decomposition can be expressed as:
$\mathcal{X}\approx\sum_{r=1}^{R}U_{1,r}^{(D_{1})}\circ U_{2,r}^{(D_{2})}\circ\dots\circ U_{n,r}^{(D_{n})}.$ (1)
In this equation, $U_{1,r}^{(D_{1})}\circ U_{2,r}^{(D_{2})}\circ\dots\circ U_{n,r}^{(D_{n})}$ represents the $r$-th rank-one tensor, where the symbol $\circ$ signifies the outer product of vectors, and $R$ denotes the rank. Specifically, $U_{n,r}^{(D_{1})}$ represents a vector corresponding to the $n$-th dimension in the $r$-th rank-one tensor, with a dimensionality of $D_{1}$.
Rethinking previous attention modules [62; 70; 4] from the perspective of CP decomposition, we can find that they are all rank-one methods. For the sake of clarity, we denote an all-one vector with a dimensionality of $D$ as $I^{(D)}$. Specifically, for TA-SNN [62], the temporal attention module for SNNs yields the attention map $\mathcal{A}_{\text{TA-SNN}}$ for the input tensor $\mathcal{X}$ as follows:
$\mathcal{A}_{\text{TA-SNN}}=I^{(HW)}_{S}\circ I^{(C)}\circ\mathcal{F}_{T}^{(T)}(\mathcal{X}).$ (2)
Here, $\mathcal{F}_{T}^{(T)}(\mathcal{X})$ represents a function that generates the temporal attention vector with a dimensionality of $T$. The same mathematical form can also be used to reconfigure the TCJA [70] and SCTFA [4] method. For TCJA, the attention map $\mathcal{A}_{\text{TCJA}}$ can be written as:
$\mathcal{A}_{\text{TCJA}}=I^{(HW)}_{S}\circ\mathcal{F}^{(C)}_{C}(\mathcal{X})\circ\mathcal{F}^{(T)}_{T}(\mathcal{X}).$ (3)
For SCTFA, the attention map $\mathcal{A}_{\text{SCTFA}}$ is:
$\mathcal{A}_{\text{SCTFA}}=\mathcal{F}^{(HW)}_{S}(\mathcal{X})\circ\mathcal{F}^{(C)}_{C}(\mathcal{X})\circ\mathcal{F}^{(T)}_{T}(\mathcal{X}).$ (4)
By comparing Equation 1 and Equation 2, 3, 4, we can easily see that the $R$ of Equation 2, 3, 4 is fixed at 1, making them special cases of Equation 1. This motivates us to design an attention module that the rank of the attention map is not fixed at 1, allowing us to control the attention effect by adjusting the rank of the attention map; namely, we wish our proposed module could be written in a form like:
$\mathcal{A}_{PFA}=\sum_{r=1}^{R}U^{(HW)}_{s,r}\circ U^{(C)}_{c,r}\circ U^{(T)}_{t,r}.$ (5)
Through this way, we can use $R$ to control certain properties of the attention map, such as low rank and the degree of compression of the original tensor. From the perspective of tensor decomposition, previous methods can be seen as special cases of our method when $R=1$. However, due to the absence of summation terms in previous methods, adjustments tailored to the specificity of the dataset cannot be made, lacking a certain degree of flexibility. The flexibility mentioned is crucial, as demonstrated by the experiments presented in Section 4.4. The results indicate that the optimal choice for $R$ varies across different datasets.